Find the current dividend on a stock, given that the required return is 9 percent, the dividend growth rate is 6 percent, and the stock price is $50 per share.
Answer #1

P0 = D0(1 + g) / (r - g)

$50 = D0(1.06) / (0.09 - 0.06)

$50 = D0(1.06) / 0.03

D0 = ($50 × 0.03) / 1.06

D0 = $1.42